# Lossy Compression with Data, Perception, and Classification Constraints

### v1

@@ V Experimental Results
For the reconstruction loss, we use MSE $\mathbb{E}(||x-\hat{x}||^{2})$ serving as the traditional distortion loss. For the perception loss, we employ the Wasserstein-1 loss to measure the distance between distributions. For the classification loss, we use the cross-entropy loss $\text{CE}(s,\mathbf{\hat{s}})$ which is shown to be an upper bound of the conditional entropy constraint[21].
@@ V-1 RDC tradeoff
To illustrate the relationship within RDC, we set the parameters in the loss function (4) as $\lambda_{p}=0$ and train the model with a series of $\lambda_{d}$ and $\lambda_{c}$. For the MNIST dataset, rates are controlled by $R=\text{dim}\times\log_{2}(L)$ with (dim, $L$) pairs including $(3,3)$, $(3,4)$, $(4,4)$, $(5,4)$. For points with the same color (i.e., trained with the same $R$), the results show that higher classification accuracy often requires sacrificing distortion, which coincides with our theoretical results. It is also observed that $D-C$ curves shift to the left-up when $R$ increases, achieving better distortion and higher classification accuracy simultaneously.
V-2 RPC tradeoff
@@
Similarly, by setting the parameters in the loss function (4) as $\lambda_{d}=0$, we illustrate RPC relationship as shown in Fig. 5(b). The figure shows that points with the same color (indicating the same rate $R$) are about parallel to the perception-axis, which coincides with our theoretical result that the rate only depends on the classification constraint and there is no tradeoff with perception constraint. When $R$ increases, we can obtain higher classification accuracy but the increment above $R\geq 8$ is marginal. We can see that under each level of accuracy (or rate), the potential of zero perception loss remains unaffected.
@@
Figure 5: Rate-distortion-classification tradeoffs in MNIST dataset. The loss is computed by setting $\lambda_{p}=0$ and varying $\lambda_{d}$ and $\lambda_{c}$.
VI Conclusion
In this paper, we considered the lossy compression problem under different constraints, including traditional distortion measure, perceptual quality characterized by divergence between distributions, and classification accuracy characterized by conditional entropy. We investigated two open problems: the information rate-distortion-classification (RDC) function and the information rate-perception-classification (RPC) function, deriving the closed-form optimal rates for both binary and Gaussian sources. Experiments on the MNIST dataset validate the theoretical results, offering insights on developing robust multi-task deep learning models and balancing diverse objectives across various domains.
@@ Appendix A Proof of Theorem 1
Recall the following information rate-distortion classification (RDC) function:
$R(D,C)=$ $\min_{p_{\hat{X}|X}}I(X;\hat{X})$ (2)
s.t. $\mathbb{E}[\Delta(X,\hat{X})]\leq D,$ (2a)
$H(S|\hat{X})\leq C,$ (2b)
where $S$ is a classification variable.
Consider a Bernoulli source $X$ and a classification variable $S$ with the binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$).
@@
Note that by data-processing inequality[15, Theorem 2.8.1] and $S-X-\hat{X}$, we have
@@
We first find a lower bound of $I(X,\hat{X})$ leveraging the constraints and then show that it is achievable. By Mrs. Gerber’s Lemma proved in [16], when $H(S|\hat{X})\leq C$, we have $H(X|\hat{X})\leq H(C_{1})$, where
$C_{1}\triangleq\frac{H^{-1}(C)-p_{1}}{1-2p_{1}},$ (5)
and $H^{-1}:[0,1]\rightarrow[0,\frac{1}{2}]$ denotes the inverse function of Shannon entropy for probability less than $\frac{1}{2}$. Note that $C_{1}=\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}\leq\frac{\frac{1}{2}-p_{1}}{1-2p_{1}}=\frac{1}{2}$. Therefore, we have
@@
$\mathbb{E}[\Delta(X,\hat{X})]$ $=P(X=0,\hat{X}=1)+P(X=1,\hat{X}=0)$
$=P(X\neq\hat{X})=P(X\oplus\hat{X}=1)\leq D,$
@@
$=H(b)-H(\hat{X}\oplus X|\hat{X})$
$\geq H(b)-H(\hat{X}\oplus X)$
$\geq H(b)-H(D).$ (7)
Since $D\leq C_{1}\leq\frac{1}{2}$, we have $H(D)\leq H(C_{1})$. Combining (6) and (7), we have the lower bound of rate $I(X;\hat{X})\geq H(b)-H(D)$. This lower bound can be achieved by choosing $(X,\hat{X})$ to have the joint distribution given by BSC with $X=\hat{X}\oplus S_{2}$ where $S_{2}\sim\text{Bern}(D)$, since $I(X,\hat{X})=H(X)-H(X|\hat{X})=H(b)-H(D)$. Here the classification constraint is also satisfied: By $D<C_{1}=\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}$, we have $p_{1}*D_{0}<H^{-1}(C)$ where $p_{1}*D_{0}=p_{1}(1-D_{0})+D(1-p_{1})$, so $H(S|\hat{X})=H(p_{1}*D_{0})<C$. Thus, $R(D,C)=H(b)-H(D)$ in this case.
@@
In summary, when $C\leq H(S_{1})$, the problem is feasible; otherwise the rate-distortion-classification function for the binary case is
$R(D,C)=\begin{cases}H(b)-H(C_{1})\qquad\text{ for }D\geq C_{1}\text{ and }C_{1}\leq b,\\ H(b)-H(D)\qquad\text{ for }D<C_{1}\text{ and }D\leq b,\\ 0\qquad\text{ for }\min\{D,C_{1}\}>b.\end{cases}$
@@ Appendix B Proof of Theorem 2
Consider the RDC problem as below,
$R(D,C)=$ $\min_{p_{\hat{X}|X}}I(X;\hat{X})$
s.t. $\mathbb{E}[(X-\hat{X})^{2}]\leq D,$
$H(S|\hat{X})\leq C,$
where $(X,S)$ are jointly multivariate Gaussian variable with covariance $\text{Cov}(X,S)=\theta_{1}$. Similar to [9, 17], the optimal reconstruction $\hat{X}$ should also be jointly Gaussian with $X$, thus the parameters to be optimized are expectation $\mu_{\hat{x}}$, variance $\sigma_{\hat{x}}^{2}$, and covariance $\text{Cov}(X,\hat{X})=\theta_{2}$.
By utilizing the formula of differential entropy and mutual information of the (joint) Gaussian variables[15, Chapter 8], the RDC problem can be represented by
@@
The second observation is that when $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$, the classification constraint is infeasible. To make $-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$ in (8) meaningful, we have $01-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}}>$ 0, i.e.,
$\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}<\sigma_{x}^{2}.$
Then, the mutual information between $S$ and $\hat{X}$ is upper bounded by
$I(S,\hat{X})$ $=-\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{4}}\cdot\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}})$
$\leq-\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}}),$
@@
In the case of (8b) not active, by Shannon’s classic rate-distortion theory for a Gaussian source with squared-error distortion[15, Chapter 10], the optimal rate is $\frac{1}{2}\log\frac{\sigma_{s}^{2}}{D}$ for $D\leq\sigma_{x}^{2}$, achieved by setting $\sigma_{\hat{x}}^{2}=\sigma_{x}^{2}-D$ and $\theta_{2}=\sigma_{x}^{2}-D$. For the classification constraint (8b), it is inactive if
$I(S,\hat{X})$ $=-\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{4}}\cdot\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}})$
$=-\frac{1}{2}\log(1-\frac{\theta_{1}^{2}(\sigma_{x}^{2}-D)}{\sigma_{s}^{2}\sigma_{x}^{4}})\geq h(S)-C,$
which is equivalent to
$D\leq\sigma_{x}^{2}-\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C}).$
@@
$I(X,\!\hat{X})\!=\!-\!\frac{1}{2}\!\log(1\!\!-\!\!\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})\!\geq\!-\!\frac{1}{2}\!\log(1\!-\!\!\frac{\sigma_{s}^{2}\sigma_{x}^{2}}{\theta_{1}^{2}}\!(1\!-\!e^{\!-2h(\!S\!)+2C})),$
with equality holds at $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}=\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{2C-2h(S)})$. Without loss of optimality, we choose $\sigma_{\hat{x}}^{2}=\theta_{2}=\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})$ and substitute it to the distortion function:
@@
When $C>h(S)$ and $D>\sigma_{x}^{2}$, we can simply choose $\sigma_{\hat{x}}=0$ (i.e., $\hat{X}$ is a constant). Then we have the rate $I(X,\hat{X})=0$, and all the constraints are satisfied:
$\mathbb{E}[(X-\hat{X})^{2}]=\sigma_{x}^{2}<D,$
$h(S|\hat{X})=h(S)<C.$
In summary, if $C<\frac{1}{2}\log(1-\frac{\theta_{1}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$, the problem will be infeasible; otherwise, the rate-distortion-classification function with MSE distortion is given by

### v2

@@ V-B1 RDC tradeoff
To illustrate the relationship within rate-distortion-classification, we set the parameters in the loss function (4) as $\lambda_{p}=0$ and train the model with a series of $\lambda_{d}$ and $\lambda_{c}$. For the MNIST dataset, rates are controlled by $R=\text{dim}\times\log_{2}(L)$ with (dim, $L$) pairs including $(3,3)$, $(3,4)$, $(4,4)$, $(5,4)$. As shown in Fig. 9(a), each point represents a encoder-decoder pair trained with a specific combination of $R$, $\lambda_{d}$, and $\lambda_{c}$. For points with the same color (i.e., trained with the same $R$), the results show that higher classification accuracy often requires sacrificing distortion, which coincides with our theoretical results. It is also observed that $D-C$ curves shift to the left-up when $R$ increases, achieving better distortion and higher classification accuracy simultaneously.
V-B2 RPC relationship
@@
By setting the parameters in the loss function (4) as $\lambda_{d}=0$ and training the model with a series of $\lambda_{p}$ and $\lambda_{c}$ under different rate $R$, we illustrate rate-perception-classification relationship as shown in Fig. 9(b). The figure shows that points with the same color (indicating the same rate $R$) are about parallel to the perception-axis, which coincides with our theoretical result that the rate only depends on the classification constraint and there is no tradeoff with perception constraint. When $R$ increases, we can obtain higher classification accuracy but the increment above $R\geq 8$ is marginal. We can see that under each level of accuracy (or rate), the potential of zero perception loss remains unaffected.
@@
Figure 9: RDC and RPC tradeoffs in MNIST dataset. (a) For the RDC tradeoff, the loss is computed by setting $\lambda_{p}=0$ and varying $\lambda_{d}$ and $\lambda_{c}$. (b) For the RPC case, The loss is computed by setting $\lambda_{d}=0$ and varying $\lambda_{p}$ and $\lambda_{c}$.
Figure 10: Rate-distortion-classification tradeoffs along different levels of perception quality for (a) MNIST, (b) SVHN datasets. We draw lines to connect the points with approximately the same W1-loss.
Figure 11: Rate-perception-classification tradeoffs given different levels of distortion for (a) MNIST, (b) SVHN datasets. We draw lines to connect the points with approximately the same MSE.
V-B3 RDC given P and RPC given D
To jointly consider distortion, perception and classification losses, we train the model with $\lambda_{d}=1$ and varying values of $\lambda_{p}$ and $\lambda_{c}$.
In Fig. 10(a) and Fig. 10(b), we show the RDP tradeoff given a certain level of perception loss on the MNIST and SVHN datasets respectively, where we connect the points with approximately the same W1-loss. For the MNIST dataset, rates are set to be $R=4,75,6,8,10$ respectively, while for the SVHN dataset, rates $R=\text{dim}\times\log_{2}(L)$ are set with (dim, $L$) pairs $(10,8)$, $(15,8)$, $(20,8)$.
Fig. 11(a) and Fig. 11(b) depict the RPC tradeoff given different levels of distortion, with the color representing the value of MSE. We draw lines to connect the points with approximately the same MSE under each rate, meaning each curve represents the relationship between classification accuracy and perception loss under a specific distortion level (e.g., 0.070, 0.063, 0.055, 0.048 for the MNIST dataset). It is observed that by jointly considering distortion loss and connecting points with the same level of MSE, the tradeoff between perception and classification emerges, which aligns with the results presented in Section IV-B.
Figure 12: Visual reconstructions for MNIST and SVHN datasets of the decoder with different $\lambda_{c}$ given $\lambda_{p}=0.015$ and $R=4.75$ (for MINST) and $\lambda_{p}=0.00125$ and $R=60$ (for SVHN) respectively.
V-B4 Improvement on the accuracy
From the experimental results, we observe that by taking the classification loss into account in our approach, sacrificing a small amount of distortion or perception quality can lead to a great improvement in classification accuracy compared to the RDP function[8, 11]. As shown in Table II, for the MNIST dataset, we set $\lambda_{d}=1$, $\lambda_{p}=0.015$ and adjust the value of $\lambda_{c}$ which controls the weight of cross-entropy loss. When $\lambda_{c}=0$, i.e., the classification is not considered, the accuracy is only $62\%$. However, as we raise $\lambda_{c}$ to 0.006, the accuracy jumps to $90\%$ with $2\%$ increase on MSE and $1\%$ increase on W1-loss. Similar results can be obtained for the SVHN dataset. As shown in Table II, when $\lambda_{c}$ increases from 0 to 0.0015, a 17% improvement in accuracy is observed, with a cost of 0.44 W1-loss and 0.001 MSE. Meanwhile, by giving classification a much larger weight ($\lambda_{c}=0.05$) as shown in the last row of Table II, we can achieve a high accuracy of approximately 0.87, approaching the accuracy tested on the original dataset using the same pre-trained classifier (approximately 0.91).
TABLE II: Comparison of distortion, perception and accuracy quality on different $\lambda_{c}$ for MNIST dataset ($\lambda_{d}=1$, $\lambda_{p}=0.015$, $R=4.75$) and SVHN dataset ($\lambda_{d}=1$, $\lambda_{p}=0.00125$ and $R=60$).
Datasets $\lambda_{c}$ Accuracy MSE W1-loss
MNIST 0 0.6218 0.0732 1.2255
0.0025 0.7612 0.0736 1.2412
0.006 0.9061 0.0747 1.2889
0.01 0.9307 0.0753 1.4128
SVHN 0 0.5744 0.0061 1.1993
0.001 0.7189 0.0068 2.2892
0.0015 0.7441 0.0070 2.4351
0.05 0.8730 0.0151 5.0038
Significant improvements can also be observed visually when incorporating the classification constraint. As shown in Fig. 12, when $\lambda_{c}=0$, numerous misclassified samples are present in the images, and it is very common for a input “4” to be reconstructed as a “9” in MNIST (or get a “8” given a input “2” and “3” in SVHN), as marked by red circles. When the weight of classification increases, such problems disappear and there is no dramatic change on the perceptual quality. However, when the weight of cross entropy loss is set too high (as $\lambda_{c}=0.05$), the distortion and perceptual quality will be degraded dramatically.
V-C Vanishment of Tradeoffs When Rate Increases
According to the results in Section IV, when there is no constraint on the rate in the lossy compression framework, or equivalently, when the noise power $\sigma_{N}^{2}$ is zero, perfect reconstruction is possible and tradeoffs should no longer exist. In this experiment, we set the rate $R$ as $125442\times 28\times 28\times\log 256=$ 12544 (i.e., twice the size of the upper bound of an original image from the MNIST dataset) and train the network using the same weights of $\lambda_{d}$, $\lambda_{p}$ and $\lambda_{c}$ as in the case of small rates.
From Fig. 13 and Fig. 14, we observe that as the rate increases, both the $(D,C)$ and $(P,C)$ points become more concentrated. By setting a sufficiently high rate (e.g., $R=12544$ bits), the tradeoffs in RDC and RPC given D are almost gone. In this scenario, the classification accuracies consistently remain at 0.975, regardless of the MSE ranging from 0.006 to 0.014, and the W-1 loss remains consistently below 0.5.
Figure 13: The vanishment of RDC tradeoff given P when the rate $R$ increases for MNIST dataset.
Figure 14: The vanishment of RPC tradeoff given D when the rate $R$ increase for MNIST dataset.
VI Conclusions
In this paper, we integrated data reconstruction tasks and generative tasks into the IB principle, focusing on the lossy compression problem with task-oriented constraints. We first investigated the rate-distortion-classification (RDC) and the rate-perception-classification (RPC) tradeoffs and derived the closed-form expressions for the RDC and RPC functions in binary and scalar Gaussian cases. These new results complement the IB principle and RDP tradeoff. Meanwhile, our analysis of lossy compression and signal restoration frameworks reveals that the presence of source noise, which can be interpreted as a specific level of distortion in lossy compression, plays a decisive role in the existence of tradeoffs. Furthermore, experiment results on a DL-based image compression justify our theoretical results, providing insights and guidelines on extracting task-relevant information to fulfill diverse objectives.
@@ Appendix A Proof of Theorem 1
Consider the RDC function in (2) with Hamming distance for a Bernoulli source $X$ and a classification variable $S$ with the binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$).
@@
Note that by data-processing inequality[17, Theorem 2.8.1] and $S-X-\hat{X}$, we have
@@
We first find a lower bound of $I(X,\hat{X})$ leveraging the constraints and then show that it is achievable. By Mrs. Gerber’s Lemma proved in [18], when $H(S|\hat{X})\leq C$, we have $H(X|\hat{X})\leq H(C_{1})$, where $C_{1}\triangleq\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}$ and $H^{-1}:[0,1]\rightarrow[0,\frac{1}{2}]$ denotes the inverse function of Shannon entropy for probability less than $\frac{1}{2}$. Note that $C_{1}=\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}\leq\frac{\frac{1}{2}-p_{1}}{1-2p_{1}}=\frac{1}{2}$. Therefore, we have
@@
$\mathbb{E}[\Delta(X,\hat{X})]$ $=P(X\neq\hat{X})=P(X\oplus\hat{X}=1)\leq D,$
@@
$\geq H(b)-H(\hat{X}\oplus X)\geq H(b)-H(D).$ (6)
Since $D\leq C_{1}\leq\frac{1}{2}$, we have $H(D)\leq H(C_{1})$. Combining (5) and (6), we have the lower bound of rate $I(X;\hat{X})\geq H(b)-H(D)$. This lower bound can be achieved by choosing $(X,\hat{X})$ to have the joint distribution given by BSC with $X=\hat{X}\oplus S_{2}$ where $S_{2}\sim\text{Bern}(D)$, since $I(X,\hat{X})=H(X)-H(X|\hat{X})=H(b)-H(D)$. Here the classification constraint is also satisfied: By $D<C_{1}=\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}$, we have $p_{1}*D<H^{-1}(C)$ where $p_{1}*D=p_{1}(1-D)+D(1-p_{1})$, so $H(S|\hat{X})=H(p_{1}*D)<C$. Thus, $R(D,C)=H(b)-H(D)$.
@@
In summary, when $C\leq H(S_{1})$, the problem is feasible; otherwise the RDC function for the binary case is
$R(D,C)=\begin{cases}H(b)-H(C_{1})~{}\text{ for }D\geq C_{1}\text{ and }C_{1}\leq b,\\ H(b)-H(D)~{}\text{ for }D<C_{1}\text{ and }D\leq b,\\ 0\qquad\text{ for }\min\{D,C_{1}\}>b.\end{cases}$
@@ Appendix B Proof of Theorem 2
Consider the RDC problem as in (2) with MSE distortion, where $(X,S)$ are jointly Gaussian variable with covariance $\text{Cov}(X,S)=\theta_{1}$. Similar to [11, 19], the optimal reconstruction $\hat{X}$ should also be jointly Gaussian with $X$, and the parameters to be optimized are expectation $\mu_{\hat{x}}$, variance $\sigma_{\hat{x}}^{2}$, and covariance $\text{Cov}(X,\hat{X})=\theta_{2}$.
By utilizing the formula of differential entropy and mutual information of the (joint) Gaussian variables[17, Chapter 8], the RDC problem can be represented by
@@
The second observation is that when $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$, the classification constraint is infeasible. To make $-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$ in (7) meaningful, we have $01-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}}>$ 0, i.e., $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}<\sigma_{x}^{2}$. Then, the mutual information between $S$ and $\hat{X}$ is upper bounded by
$I(S,\hat{X})=-\frac{1}{2}\log(1\!-\!\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{4}}\!\cdot\!\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}})\!\leq\!-\frac{1}{2}\log(1\!-\!\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}}),$
@@
In the case of (7b) not active, by Shannon’s classic rate-distortion theory for a Gaussian source with squared-error distortion[17, Chapter 10], the optimal rate is $\frac{1}{2}\log\frac{\sigma_{s}^{2}}{D}$ for $D\leq\sigma_{x}^{2}$, achieved by setting $\sigma_{\hat{x}}^{2}=\sigma_{x}^{2}-D$ and $\theta_{2}=\sigma_{x}^{2}-D$. For the classification constraint (7b), it is inactive if
$I(S,\hat{X})$ $=-\frac{1}{2}\log(1-\frac{\theta_{1}^{2}(\sigma_{x}^{2}-D)}{\sigma_{s}^{2}\sigma_{x}^{4}})\geq h(S)-C,$
which is equivalent to $D\leq\sigma_{x}^{2}-\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})$.
@@
$I(X,\!\hat{X})\!\!\geq\!-\!\frac{1}{2}\!\log(1\!-\!\!\frac{\sigma_{s}^{2}\sigma_{x}^{2}}{\theta_{1}^{2}}\!(1\!-\!e^{\!-2h(\!S\!)+2C})),$
with equality holds at $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}=\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{2C-2h(S)})$. Without loss of optimality, by choosing $\sigma_{\hat{x}}^{2}=\theta_{2}=\sigma_{s}^{2}\sigma_{x}^{4}(1-e^{-2h(S)+2C})/{\theta_{1}^{2}}$ and substituting it to the distortion function, we can obtain
@@
When $C>h(S)$ and $D>\sigma_{x}^{2}$, we can simply choose $\sigma_{\hat{x}}=0$ (i.e., $\hat{X}$ is a constant). Then we have the rate $I(X,\hat{X})=0$, and all the constraints are satisfied: $\mathbb{E}[(X-\hat{X})^{2}]=\sigma_{x}^{2}<D$ and $h(S|\hat{X})=h(S)<C$. In summary, if $C<\frac{1}{2}\log(1-\frac{\theta_{1}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$, the problem will be infeasible; otherwise, the rate-distortion-classification function with MSE distortion is given by